# Not Everyone Wins with LLMs: Behavioral Patterns and Pedagogical Implications in AI-assisted Data Analysis

### v1

@@ 6.2. Analysis results: AI usage change across steps and expertise
We believe all of these expertise gaps might be closed through training on AI use skills, e.g., when to query AI and with what prompt. For example, a sequence of ai_breakdown_intent $\rightarrow$ ai_generate_code might help a novice obtain more easily digestible code in the first place; if followed by “further simplify your explanation like I’m five”, their chance of correctly generating and understanding AI code can be improved. Similarly, students can start by asking for ai_suggest_steps_or_plans to familiarize themselves with the task, or ai_explain_concepts to understand domain-specific keywords. However, our analysis shows that much of the student grade and effective AI usage still comes down to tech expertise. In the next section, we explore what we should we do better in teaching AI use skills.
7. RQ3: What and how can we train students on AI usage?
@@
The findings in RQ2 suggest that technical experts also use AI more strategically and effectively than novices. Here, we ask: How can we train students to use AI more effectively? What are some skills and learning opportunities for training better AI users?
7.1. The effect of in-class demonstration and extended time-on-task
While we did not experiment with intervention mechanisms for training AI use skills (since the core learning objective was on computational data analysis), we did collect some preliminary signals on what AI use skills might be more easily acquirable than others via just in-class demonstrations and extended time-on-task.
Figure 8. Appropriate AI use ratio pre- to post-instructions. Ratio = $used/(used+missed)$.
Method.
Between HW0 and HW1, the instructor provided a 90-minute live-coding demonstration of the HW0 data cleaning task. The demonstration showcased examples of AI use behaviors in the codebook (Table 3), with explanations of when and why each behavior could be useful. For example, the instructor explained “If you find unfamiliar concepts, you can actually ask Gemini to get you started. We can ask, What are some key steps in Data cleaning?” while typing and sending the actual question and walking through the Gemini’s answer. The demo also covered all the task environment space (e.g., overview of Colab as a product) as well as the AI context space (e.g., the code cell generation has access to the entire notebook, while the chatbot sidebar does not have access to code cell outputs). Students could also re-watch the recording asynchronously. Despite the limitations of lecturing (e.g., less cognitively engaging than active learning and may create an illusion of fluency without fostering durable understanding (Chi and Wylie, 2014; Deslauriers et al., 2019)), this is a lightweight instruction baseline.
@@ 7.1. The effect of in-class demonstration and extended time-on-task
In addition to instructions on AI use (and regular lectures on programming and data science domain knowledge), students moved from HW0’s 15-minute time-limited tasks to HW1 and HW2’s bi-weekly assignments (Table 1). Therefore, when we analyze students’ behavioral changes pre- to post-demo, any improvement is a result of both extended time on task and the instruction, and students’ room for improvement builds on top of gains already afforded by extended practice time and baseline instructional exposure.
Figure 9. The distribution of the last successful step in an episode, before and after class demonstration.
Result.
As shown in Figure 9, students were able to sustain progress further along the problem-solving process at post-demo. In HW0, many episodes stalled at early stages (examples of student reported challenges in Table 7; Appendix A) or ended without any successful step (already failed at intent, denoted by “none”). After provided with baseline instructions and ample time, more episodes advanced to later steps of understanding and assessment.
We also computed an appropriate AI use ratio ($used/(used+missed)$) to measure whether students capitalized on available AI opportunities. As shown in Figure 8, appropriate AI use ratios increased post-demo in many behaviors. While not fully mastered or statistically compared, there is a notable progress in students’ prompt quality behaviors by an average of 30% in AI use ratio. Students also showed modest gains in asking ai_explain_concepts. However, evaluation behaviors remained a bottleneck (e.g., post-demo ratios of using AI to critique, asking for code or bug explanations are all ¡ 50%). This suggests a shared hesitation to request or process complex evaluative information and use AI to deepen understanding. These observations hint towards essential but under-taught AI use skills, such as the disposition to persist through trial-and-error, read and interpret long outputs, and refine AI’s outputs (more in Section 7.2).
Note that some AI use behaviors like provide_context_in_prompt are much more frequent (18.06%) than others (Figure 4A), like ai_improve_prompt (0.17%), decompose_task_in_prompt (0.31%), ai_edit_partial_code (0.57%), and ai_breakdown_intent (0.65%). The appropriate AI use ratios of infrequent categories can be less interpretable and easily exaggerated by one or two log instances.
@@
7.2. What should we teach on AI use skills
As shown in Figure 8, we observe much room for improvement in AI use skills.To produce actionable insights for instructors, we identified lingering AI use challenges that are not solved by simple classroom demonstrations (using both survey responses and log examples), and summarized concrete learning objectives.
Method
@@ 7.2. What should we teach on AI use skills
From five post-surveys’ answers to the questions “What were the challenges for you during the tasks when you used GenAI” and “What do you think could help you further improve your performance with GenAI on these tasks?”, we derived learning objectives for AI use skills (Table 4). We adopt the knowledge dimensions taxonomy to categorize AI use skills into conceptual, procedural, dispositional, and metacognitive knowledge (Krathwohl, 2002; Wiggins and McTighe, 2005). Using a similar thematic analysis procedure as described in Section 6.1.2, we iteratively constructed the codebook guided by an open coding process (Strauss and Corbin, 1998), and we validated the codes with our log analysis (Table 6).
Result
@@
In short, effective AI use is not simply reduced to “good prompts” — AI challenges that students encountered often stem from a lack of various AI use skills. As shown in Table 4, we revealed necessary conceptual knowledge (what to know about AI), procedural skills (how to use AI), metacognitive skills (when to apply AI), and dispositional skills (keep engaging with AI). [^5] There can be affective and motivational learning goals, such as self-efficacy and interest to explore. However, these goals are not manifested in our behavioral log data, and we discuss more about possible extensions in future work (Section 8.2).
@@
Table 5. Comparison of expert vs. novice behaviors in resolving KeyError: Column not found. Experts identify missing context (e.g., actual column names) and fix issues using grounded prompts or direct code edits. Novices often rely on vague prompts or trial-and-error, leading to unresolved errors.
Edit on Expert Behavior Novice Behavior
@@
 Uses “Explain Error” → LLM suggests fixing to Price
 Applies suggestion → resolves error
@@
Differentiate Knowledge Dimensions
Below, we provide some behavioral examples to help illustrate each AI use dimension. However, from our behavioral log data, we could not always identify which knowledge dimension exactly applies. For example, a student may know a particular way of using AI already; however, they may not be able to retrieve it properly at the time they need to use it (available but not accessible knowledge (Tulving and Pearlstone, 1966)). Especially when a student did not use AI, we may not conclude if it was because they did not notice the appropriate time to apply this knowledge (metacognitive), did not have this knowledge stored (conceptual), failed to execute properly (procedural), or lacked the persistence to keep trying (dispositional). Future work may design more targeted assessments corresponding to each type of knowledge in order to differentiate them (Section 8.2).
Metacognitive.
Students often struggled with deciding when to involve the AI and what to delegate. As we analyzed HW1&2 logs (after class instructions), we found multiple instances of students getting stuck in a loop of explain_error and generate_code without re-evaluating their approach (such as the novice’s code edit sequence Table 5). Specifically, these failures stem from repeated KeyError exceptions caused by the AI generating code with nonexistent columns — as Gemini in Colab was not able to retrieve the correct column names from the dataframe, the new code without augmented context kept hallucinating nonexistent columns.
This pattern could illustrate multiple missing AI use skills, such as a lack of metacognitive monitoring: the ability to recognize when a strategy is unproductive and to adjust. The students could avoid delegating AI to do the task that it has demonstrated unreliability and providing the correct column name themselves (such as the expert’s code-edit sequence Table 5), or alternatively ask AI to breakdown the debugging procedure and walk them through a step-by-step process to figure out what is wrong, or add the exact column name as context_in_prompt (such as the expert’s prompt-edit sequence Table 5). Effective metacognitive use would involve pausing to decide, for example, whether to ask the AI for a structured plan (e.g., “list all columns in my dataset”) before persisting with trial-and-error.
@@
Many failures could stem from the lack of conceptual knowledge regarding the AI’s task space (what it can do) and context space (what information it can access). For task space, most showed improvement in appropriate AI use (Figure 8); however, students may not learn everything they can ask AI to do with simple demonstrations. For instance, the instructor covered using AI to rewrite and improve_prompt, which appropriate use increased from 0% to 100% in post-demo (despite low frequency; Figure 4). Nevertheless, some students still did not realize this use until the later half of the semester: “now I found a smooth way to utilize GenAI – ask it to also re-write my prompt before I use the prompt to generate the code.” (HW3 post-survey). If the demonstration did not covered some AI use, students may develop a wrong assumption of the AI task space. For example, the instructor did not cover ai_edit_partial_code during the demo, and a student demonstrated struggles caused by the lack of this AI use knowledge in their post-survey: “you can’t fine-tune parts of the result; instead, you have to regenerate the whole thing, and sometimes, parts that were correct before end up wrong.”
@@
Procedural skills concern how to best interact with the AI. Just knowing what AI needs is not enough, as one student noted, “Most of the time, it was difficult to feed context because there were multiple sources to refer to” (post-class survey), reflecting the challenge of providing the appropriate context in a prompt (procedural) even with the conceptual knowledge of AI context space. Similarly, one may know they should break down a complex goal without being able to accomplish it, and end up settling for downgraded goals: “I didn’t really know how to write the prompt for the type of complex data analyses to generate the charts that I wanted, so I had to keep to simple analyses and charts.” In the demo, the instructor exemplified good prompting strategies, such as writing clear_instruction, providing_context, and decomposing_task_in_prompt. However, the post-class behaviors still leave 20% of missing AI use opportunities as room for improvement (Figure 8).

### v2

@@ 6. RQ2: How do experience differences translate to AI usage behaviors?
We believe these experience gaps might be closed by training on AI use skills. For example, a sequence of ai_breakdown_intent $\rightarrow$ ai_generate_code might help a novice obtain more easily digestible code; if followed by “further simplify your explanation”, their chance of correctly understanding AI code can be improved. Similarly, students can ask for ai_suggest_steps_or_plans to familiarize themselves with the task, or ai_explain_concepts to understand domain-specific keywords. However, our analysis shows that much of the student grade and effective AI usage still comes down to technical experience. In the next section, we explore what we should we do better in teaching AI use skills.
7. RQ3: What can we train students on AI usage?
@@
The findings in RQ2 suggest that technically experienced students use AI more strategically and effectively than novices. Here, we ask: What are some skills and how to teach better AI users?
@@
By comparing student behaviors pre- to post-demo (Section 3.3), we collected preliminary signals on what AI use skills might be more easily acquirable than others. Note that in addition to instructions on AI use (and regular lectures on programming and data science knowledge), students moved from HW0’s 15-minute tasks to bi-weekly assignments (Table 1). Therefore, any improvement pre- to post-demo is a result of both extended time-on-task and the instruction, and any room for improvement builds on top of gains already afforded by extended practice time and baseline instruction.
@@
Learning objectives for AI-use skills grouped by cognitive, conceptual, procedural, and dispositional dimensions, describing what students should know and how they should engage when using AI tools.
Figure 7. The distribution of the last successful step in an episode, before and after class demonstration. More students post-instructions can successfully get to last step of assess in their workflow comparing to HW0.
Bar chart showing distribution of the last successful step in each analysis episode before and after instruction. Post-instruction episodes progress further, especially to the assessment stage.
Figure 8. Appropriate AI use ratio pre- to post-instruction. Ratio = $used/(used+missed)$.
Bar charts comparing pre- and post-instruction appropriate AI use ratios across improvement, coding, explanation, evaluation, and prompt-quality behaviors. Prompt-quality behaviors show the largest gains; evaluation remains limited.
In-class demonstration and extended time-on-task improved AI use success, especially by improving prompt, but evaluation behaviors need more targeted instructions.
As shown in Fig. 7, students were able to sustain progress further along the problem-solving process at post-demo. In HW0, many episodes stalled at early stages (examples of student-reported challenges in Table 6; Appendix A) or ended without any successful step (already failed at intent, denoted by “none”). After being provided with baseline instructions and ample time, more episodes advanced to later steps in the sequence of intent $\rightarrow$ input $\rightarrow$ understand $\rightarrow$ assess (as described in Section 4.3).
We also computed an appropriate AI use ratio ($used/(used+missed)$) to measure whether students capitalized on available AI opportunities. As shown in Fig. 8, appropriate AI use ratios increased post-demo in many behaviors. While not fully mastered or statistically compared, there is a notable progress in prompt quality behaviors by an average of 30% in students’ AI use ratio. Students also showed modest gains in asking ai_explain_concepts. However, evaluation behaviors remained a bottleneck (e.g., post-demo ratios of using AI to critique, asking for code or bug explanations are all ¡ 50%). This suggests a shared hesitation to request or process complex evaluative information and use AI to deepen understanding. These observations hint towards under-taught AI use skills, such as the disposition to persist through trial-and-error, read and interpret long outputs, and refine AI’s outputs.
Note that some AI use behaviors like provide_context_in_prompt are much more frequent (18.06%) than others. The appropriate AI use ratios can be less interpretable and easily exaggerated by one or two log instances for infrequent categories like ai_improve_prompt (0.17%), decompose_task_in_prompt (0.31%), and ai_breakdown_ intent (0.65%) (details in Fig. 9 in Appendix A).
Learning objectives of AI use skills across knowledge dimensions.
@@
To produce actionable insights for instructors, we examined both survey responses and log examples to identify AI use challenges that persist beyond classroom demonstrations and to distill concrete learning objectives. From all 36 students’ five post-survey answers to the questions “What were the challenges during the tasks when you used GenAI” and “What do you think could help you further improve your performance with GenAI on these tasks?”, we derived learning objectives for AI use skills (Table 4). We adopted the knowledge dimensions taxonomy to categorize AI use skills (Krathwohl, 2002; Wiggins and McTighe, 2005). Using a thematic analysis procedure similar to Table 2, we iteratively constructed the codebook via open coding (Strauss and Corbin, 1998), and we validated the codes with examples from log data (details in Table 7, Appendix A).
Differentiate Knowledge Dimensions
@@
In short, AI challenges that students encountered often stem from a lack of various AI use skills. In Table 4, we revealed necessary conceptual knowledge (what to know about AI), procedural skills (how to use AI), metacognitive skills (when to apply AI), and dispositional skills (keep engaging with AI). [^7] There can be affective and motivational learning goals, such as self-efficacy and interest to explore. However, these goals are not manifested in our behavioral log data, and we discuss more about possible extensions in future work (Section 8.2). We provide some behavioral examples to help illustrate each AI use dimension. However, from our behavioral logs, we could not always identify which knowledge dimension exactly applies. For example, a student may know a particular way of using AI already, but they may not be able to retrieve it properly at the time they need to use it (available but not accessible (Tulving and Pearlstone, 1966)). Especially when a student did not use AI, we may not conclude if it was because they did not notice the appropriate time to apply this knowledge (metacognitive), did not have knowledge stored (conceptual), failed to execute properly (procedural), or lacked the persistence to keep trying (dispositional). Future work may design more targeted assessments corresponding to each type of knowledge (Section 8.2).
Metacognitive.
Students often struggled with deciding when to involve the AI and what to delegate. As we analyzed HW1&2 logs (after class instructions), we found multiple instances of students getting stuck in a loop of explain_error and generate_code without re-evaluating their approach (such as the novice’s code edit sequence Table 5). These failures stem from repeated KeyError exceptions — as Gemini in Colab was not able to retrieve the correct column names from the dataframe, code generation without augmented context kept hallucinating nonexistent columns.
@@
Table 5. Comparison of experienced students vs. novice behaviors in resolving KeyError: Column not found. Experienced students identify missing context (e.g., actual column names) and fix issues using grounded prompts or direct code edits. Novices often rely on vague prompts or trial-and-error, leading to unresolved errors.
Edit on Technically Experienced Behavior Novice Behavior
@@
 Uses “Explain Error” → LLM suggests fixing to Price → applies fix → error resolved
@@
Side-by-side examples comparing how experienced and novice students resolve KeyError issues. Experienced students inspect data and correct column names, while novices repeatedly use vague prompts and often fail to resolve the errors.
This pattern could illustrate multiple missing AI use skills, such as a lack of metacognitive monitoring: the ability to recognize when a strategy is unproductive and to adjust. Students could avoid delegating AI to do the task that it has demonstrated unreliability and providing the correct column name themselves (such as the experienced student’s code-edit sequence Table 5), or alternatively ask AI to breakdown the debugging procedure and walk them through step-by-step, or add the exact column name as context_in_prompt (such as the experienced student’s prompt-edit sequence Table 5). Effective metacognitive use would involve pausing to decide, for example, whether to ask the AI for a structured plan (e.g., “list all columns in my dataset”) before trial-and-error.
@@
Failures could stem from the lack of conceptual knowledge regarding the AI’s task space (what it can do) and context space (what information it can access). For task space, most showed improvement in appropriate AI use (Fig. 8); however, students may fail to learn with simple demonstrations. For instance, the instructor covered using AI to improve_prompt, which appropriate use increased from 0% to 100% in post-demo (despite low frequency of 0.17% per student). Nevertheless, some students still did not realize this use until the latter half of the semester: “now I found a smooth way to utilize GenAI – ask it to also re-write my prompt before I use the prompt to generate the code.” (HW3 post-survey). If the demonstration did not cover some AI use, students may develop a wrong assumption. For example, the instructor did not cover ai_edit_partial_code during the demo, and a student demonstrated struggles caused by the lack of this AI use knowledge in their post-survey: “you can’t fine-tune parts of the result; instead, you have to regenerate the whole thing, and sometimes, parts that were correct before end up wrong.”
@@
Just knowing what AI needs is not enough, as students still need skills to best interact with the AI. One student reflected on the challenge of providing appropriate contexts in prompts (procedural), even when knowing AI context space (conceptual): “Most of the time, it was difficult to feed context because there were multiple sources to refer to” (post-class survey). Similarly, one may know they should break down a complex goal without being able to accomplish it, and end up settling for downgraded goals: “I didn’t really know how to write the prompt for the type of complex data analyses to generate the charts that I wanted, so I had to keep to simple analyses and charts.” In the demo, the instructor exemplified good prompting strategies, such as writing clear_instruction, providing_context, and decomposing_task_in_prompt. However, post-class behaviors still leave 20% of missing AI use as room for improvement (Fig. 8).
Dispositional.
The lack of persistence may also result in early stops. For instance, when a student tried to add labels on top of a pie chart, the generated code produced errors, and simply pressing the “explain error” button did not fix the issue. The student stopped trying, accepting the basic pie chart without the label and commented “seems like it either do the label or a pie” (HW2 think-aloud). In comparison, success is shown when a student keeps trying to refine outputs until they achieve the intended goals, e.g., persisting through different AI mistakes to produce the ideal visualization:
$\rightarrow$ regenerate to make the grouped bar chart
$\rightarrow$ Change this to output one such chart per year. There should be one chart for every year in the range.
(HW2 log)
8. Discussion and Future Work
8.1. Design and Pedagogy Implications
Make AI mental models explicit.
Unlike simple chat interfaces widely studied by existing works (Chen et al., 2024a; Sawalha et al., 2024; Tassoti, 2024; Nguyen et al., 2024), AI integrated in a notebook exposes multiple affordances (code cells, outputs, error panes, menu buttons, etc.). We observed in the study that students transferred assumptions from more capable general-purpose AIs to their Colab working environment, miscalibrating what the embedded model could access or do. This produced confusion about AI task space (what to ask) and context space (what the model can “see”). In post-surveys, some students commented that they switched to other AI tools like ChatGPT because Gemini keeps producing bugs or have to prompt to exclude the exact column key every time. As shown in Table 5, a novice prompted the model “Don’t use nonexistent columns, use what you have in the datasheet” after three repeated KeyError and commented “Perplexity can do that,” assuming that Colab Gemini would have access to the imported dataframe like other systems. It took them more than five repeated failures to realize that they should adjust their mental model for the system at hand, highlighting the need for explicit conceptual models and metacognitive abilities.
Similar issues exist in the long history of research on mental models of software systems (Council and Others, 1987). On the design side, systems with GenAI assistants should make their task and context scope more legible (e.g., what files and variables are visible to GenAI, what operations are supported, etc). Teaching “what to ask this specific AI” and “what this specific AI knows” is likely getting obsolete as models and systems rapidly evolve (Ma et al., 2025). More importantly, instructions are needed for metacognitive abilities to adjust mental models to different AI system capabilities, and methods such as self-reflection may help expand AI literacy to metacognitive skills (Kumar et al., 2024).
As we found in Section 7, using AI to gain additional, evaluative information was a consistent challenge. Asking AI to explain code, bugs, or critique outputs remained limited, despite its potential benefit to improve understanding. We especially observe that students may not persist in reading convoluted output, reflecting reluctance or difficulty in engaging with evaluative or complex information. While there are one-click affordances (explain error buttons) that aim to make it easier to start evaluation, the cognitive load of seeing a lengthy error explanation often creates resistance or reduces the chance for students to learn critical evaluation skills.
8.2. Limitations and Future Work
Our analyses suggest that students’ technical experience significantly predicts performance during LLM use (similar to prior literature (Chen et al., 2024a)), whereas LLM and communication experience factors did not. Additionally, we showed that the technical experience gap was reduced by LLM use under time pressure. Other prior work also shows that LLMs can narrow gaps between novices and experts (Pickering et al., 2025), or even impede expert performance under certain conditions (Becker et al., 2025). While we did not experimentally isolate the effect for time-on-task, we provide preliminary evidence that time constraints may influence the effectiveness of LLM use. The different results of whether LLM closes the experience gap or not in bi-weekly assignments versus tasks in lecture highlight that technical experience is not the only factor shaping effective LLM use (Section 5).
Our sample included 36 students in authentic classroom contexts, which ensured ecological validity. However, it is a small sample that constrained our ability to derive conclusive subgroup analysis and isolate factors, such as undisclosed external AI use and overlapping experience dimensions. For example, Python and data science proficiency were moderately correlated in our sample, so we could not examine their independent effects on effective LLM use. Our log analyses revealed the importance of AI use strategies, which point toward the need to define and evaluate “AI use experience” more directly, as what we have attempted in Section 7. We also measured affective factors (motivation, self-efficacy) in self-reported survey items (Appendix C), but none significantly predicted performance. Experience scores in our work were operationalized as self-reported survey items, which, despite being common practice (Chen et al., 2024a; Lum et al., 2025; Chen et al., 2024b), have some known issues such as miscalibration of self-perceived experience (Kruger and Dunning, 1999). We partially validated our technical experience measure with the performance in HW0 tasks without using LLMs; however, future studies could recruit more distinct participant groups or employ more rigorous pretests. More robust performance-based assessment and controlled experiments are needed to disentangle the interplay between more nuanced factors such as domain experience, implementation experience, AI-use experience, affect, and situational factors like time pressure.
Going broader: extend to more diverse task contexts.
While the learning objectives for AI use skills (Table 4) were derived from programmatic data analysis tasks, they should extend beyond our study context. Many objectives are relevant to problem-solving and higher-order thinking skills, such as clarifying the task space of any tool, communicating requirements with decomposition, and verifying results. Future research could examine whether instruction on AI use transfers to more general competencies, and explore how to evaluate AI use experience more systematically (as self-rated LLM experience may not be appropriate).
Finally, we chose Google Colab for the class because it embeds a free, in-house Gemini, which enabled us to explore diverse behavioral AI use patterns in our research questions. Compared to prior work that analyzes linear chat logs (Lee et al., 2022), our setup in Google Colab involves multiple LLM-powered functions with different affordances and model access, resulting in more nuanced observations about when, how, and why students turn to AI support (e.g., comparing behaviors between AI-as-chat versus AI-as-inline-tool). However, our data collection relied on a custom logging script in the browser, which can be less robust than existing Jupyter notebook logging toolkits like JELAI (Valle Torre et al., 2025). For example, our approach may capture an auto-completion as users’ code edits (although infrequent), and Colab’s frequently changing interface raises concerns about long-term replicability. Future work could build on established logging toolkits to improve data quality and generalizability.
9. Conclusion
LLMs promise to broaden access to technical work, yet our study shows that not everyone benefits equally. Technical experience remained the strongest predictor of success in homework, and more technically experienced students engaged with AI more strategically than novices. Behavioral log analysis highlighted that experienced students provide clearer context, decompose tasks, and strategically use AI, while novices tend to rely on AI reactively in the face of challenges. While demonstrations and extended time helped students adopt clearer prompting practices, evaluative skills remained challenging. These findings suggest that effective AI use needs to cultivate durable competencies in multiple dimensions. We translate these insights into design and pedagogy implications for AI literacy: explicate mental models, clarify before generating, and scaffold verification. Ultimately, effective AI use is not just better prompts — it is about better problem solving.
Acknowledgements.
Thanks to all the participants for the study, and thanks to all LearnLab members who provided feedback on this work. Thanks to the National Science Foundation (award CNS-2213791, 2414915) and Google Academic Research Award for partial support of this work. Acknowledgment of LLM Use. We acknowledge the use of ChatGPT to assist with language editing, summarization, and restructuring of text during manuscript preparation. Claude was used to assist the analysis of log data; implementation details are included in Table 2, prompts and settings are attached in Appendix E. Copilot was used to assist the visualization and code writing to process and clean our dataset, and Gemini in Google Colab was used to prepare the homework materials and reference solutions.
Form
Input
Table summarizing student-reported challenges when using GenAI, such as inaccurate outputs, unclear explanations, and difficulty refining prompts.